OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

-3(x + 7) = - 12 =-3x -21 = - 12 = x = - 12 + 3 x = -9

OpenStudy (anonymous):

-3(x + 7) = -12 =-3x -21 = - 12 -3x = -12 +21 -3x = 9 x = 9/ -3 x = -3

OpenStudy (anonymous):

Sorry x = 9 / 3 x = 3